Shift report 20011103 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sat Nov 03 2001 - 07:49:53 EST

  • Next message: Apache: "Shift report 20011103 08:00-16:00"

    ------------------------------------------------------------
    Date: Sat Nov  3 07:49:53 2001, Shift: 00:00-08:00
    Supervisor: rw, with: 
    Log book#10 pages 90 - 91
    
    Shift Summary:
    ------------------------------------------------------------
    1.After three hours of machine study for improving the life time of Blue ring, 
    a marginal storage was made at 3:00am. Blue/Yellow=15.7/15.6
    ZDC rate was 350. Physics run started at 3:15. At 5:30 beam was dumped.
    Soon after QLI was announced. 6:00 QLI recovered. After one ramping failed, 
    the machine went into beam diagonostic study using 6 bunch beam.
    
    2.  run     Trigger 6    
        5641      17K           FS 12deg-1/4B, MRS 60deg-3B
        5642     100K           same
        5643     TPC Pedestal run
        5644     DVM run
        5645     Pedestal run
    
    3. For T5 mod 1 and mod 3, all three Cathods tripped off during run 5641.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sat Nov 03 2001 - 07:50:30 EST